Find the mean of the probability distribution of the random variable x, which can take only the values 1, 2, and 3. Given that P(1) = 10/33, P(2) = 1/33, and P(3) = 12/33


1
Expert's answer
2022-03-01T06:05:37-0500

Check

"\sum _ip(x_i)=10/33+1/3+12/33=1"


"mean=E(X)=\sum_ix_ip(x_i)"

"=1(10\/33)+2(1\/3)+3(12\/33)=58\/33"



The mean of the probability distribution of the random variable "x" is "58\/33."